[Date Prev][Date Next][Thread Prev][Thread Next][Search]
[HEASARC Mailing List Archives]
Compton Gamma Ray Observatory Status Report #195
Monday, August 12, 1996
Questions or comments can be sent to
Chris Shrader at the CGRO-SSC.
Guest Investigator News:
The CGRO Timeline committee met in late July to determine the
Cycle-6 Viewing Plan. The plan has now undergone constraint
checking and review and it will be posted on the CGRO SSC
WWW page shortly, as will a detailed summary of accepted
proposals and all on- and off-axis GI targets. About 95%
of the GI notification letters have been sent out and the
remainder will be sent shortly. Thanks again to all of you
who proposed for your continuued interest in the program.
The high voltage for the track imaging system of EGRET was
reactivated on July 30, 1996. It had been off since June 14,
1996 to conserve the remaining gas, since the astrophysical
objects in the field of view were judged by the proposal review
committee to be of lesser priority. Delivery of the final phase
4 data to the GRO SSC is on schedule, and delivery of the phase 5
preliminary data to the GRO SSC is also on schedule. Interaction
with guest investigators continues at a good level.
A paper entitled "EGRET High-Energy Gamma-Ray Pulsar Studies.
III A Survey" by Nel et al. appeared in the Astrophysical
Journal, 465, 898 (1996) and a paper entitled "EGRET Observations
of High-Energy Gamma Radiation from PSR B1706-44" appeared in the
Astrophysical Journal, 465, 385 (1996). Also, all of the EGRET
Instrument Science Team papers presented at the Third Compton
Symposium are now available in preprint form.
In case that you missed previous announcements, information on
EGRET is now available on the World Wide Web. You may access it
Currently, the EGRET telescope is pointed at Geminga and will
move to the Crab on August 13, 1996.
We continue to prepare for the Fourth Compton Observatory
Symposium to be held in Williamsburg VA on 27-30 April 1997. For
more information, contact Jim Kurfess (email@example.com)
or check the Symposium home page at URL
OSSE operations are normal. The instrument is working as
designed, with all subsystems in complete and full operation.
The slewing response to BATSE burst triggers was disabled on
96/213 as part of our engineering tests on the Crab in the
current viewing period. It should be restarted on 96/233.
A number of preprints and reprints of recent OSSE results have
just been mailed. Preprints are also available on the OSSE WWW
page at URL http://osse-www.nrl.navy.mil/osselib.htm.
Recent observations are listed in the following table.
View period Dates Target (owner)
524 9-23 July GX 339-4 (E. Liang)
Vela Pulsar (A. Harding)
525 23-30 July Vela Pulsar (A. Harding)
PKS 2155-304 (public)
NGC 7172 (public)
NGC 7213 (public)
526 30 July - 13 Aug Crab Pulsar (PI team)
IC 4329A (PI team)
NGC 2992 (public)
Low-level OSSE data products through viewing period 415 and
high-level data products through viewing period 220 have been
delivered to the Compton GRO Science Support Center archive. In
addition, by special request all subsequent public Cyg X-1 data
sets, both low and high level, have been delivered. Refer to the
CGRO-SSC page on the WWW (http://cossc.gsfc.nasa.gov), or contact
Tom Bridgman (firstname.lastname@example.org) for more
The COMPTEL instrument is performing well and continues routine
The Sun was within the field of view of COMPTEL on 9 July 1996,
when the active region AR-7978 produced a rare (during this
period of minimal solar activity), powerful X2.9-class flare.
COMPTEL obtained gamma-ray light curves of this event in three
operational modes of the instrument (double-scatter telescope,
and high- and low-energy burst detectors). Preliminary spectral
results show a number of nuclear emission lines, for which
detailed analysis is in progress.
A general team meeting of the collaboration was held during the
week of 23-26 July at MPE in Garching, Germany. At the recent
COSPAR meeting in Birmingham, England, members of the
collaboration presented recent research results on the topics of
pulsars (Hermsen et al.), gamma-ray bursts (Kippen et al.), and
neutron albedo measurements with COMPTEL (Morris et al., and
Weidenspointner et al.).
Lastly, though losing the 83rd Tour de France to the valiant
Dane, Bjarne Riis, the previously unbeatable Indurain achieved
some measure of vindication by winning Olympic gold in the
inaugural cycling time-trial event in Atlanta.
The following was reported in IAU circular 6432:
M. H. Finger and M. Scott, Universities Space Research
Association; T. A. Prince and B. Vaughan, California Institute of
Technology; and R. B. Wilson, Marshall Space Flight Center, NASA,
report for the Compton Observatory BATSE team: "Renewed activity
from the x-ray-transient pulsar 2S 1417-624 is being observed.
Hard x-ray pulses were first detected on June 30, with the pulsed
flux rising to 17 +/- 3 mCrab (20-50 keV) by July 8. A
frequency of 0.05694816(25) Hz on July 4.5 UT and barycentric
frequency rate of 3.35(14) x 10E-11 Hz/s were obtained from a fit
of frequency measurements from July 1-8. The source had been
undetected by BATSE since a series of six outbursts were
beginning on 1994 Aug. 26 (IAUC 6075) and ending 1995 July 8."
The following was reported in IAU circular 6436:
GRO J1655-40 AND GRS 1915+105
B. A. Harmon, C. A. Wilson, M. McCollough, S. N. Zhang, W.
Paciesas, and C. R. Robinson report for the BATSE Compton
Observatory Science Team: "The two superluminal jet sources GRO
J1655-40 and GRS 1915+105 are both being detected strongly in
x-rays by BATSE. GRO J1655-40 was first detected in late May
but has gradually increased in brightness to an average around
Crab (20-100 keV) beginning about July 1. Occasional daily
averages reach 0.9 Crab. The spectrum is similar to that of
previous outbursts, exhibiting a power law shape to 200 keV with
photon indices of -2.3 to -2.6. GRS 1915+105 has also become
active again in hard x-rays, and is undergoing a moderately
outburst (about 200 mCrab in the band 20-100 keV) which began
around July 4. This precedes the increased radio flux by about 5
days (IAUC 6432). Recent increases in hard x-ray flux (around
June 21 and July 4 onward) have been accompanied by spectral
softening. The spectrum, adequately fit by a power law, has a
spectral index varying between -2.4 and -3.4."
During the last month the following pulsed sources have been
detected by the
BATSE pulsed source monitor: Her X-1, Cen X-3, 4U 1626-67, 2S
EX 02030+375, OAO 1657-415, GX 1+4, Vela X-1, and GX 301-2.
The burst trigger is currently using count rates for events above
As of August 7 BATSE has detected 1622 gamma-ray bursts out of a
total of 5462
on-board triggers in 1933 days of operation. There have been 768
to solar flares, 10 due to SGR events, 55 due to terrestrial
and 1477 due to the bursting pulsar GRO J1744-28.